# Latent learning: episodic memory complements parametric learning by enabling flexible reuse of experiences

### v1

@@ 4.3 Semantic structure dataset
This dataset is adapted from the semantic structure benchmark of Lampinen et al. (2025); however, as we are training models from scratch rather than finetuning, we make a few modifications. We generate data using the same generating process from a larger structure of 1,100 entities, sample a larger number of documents (11,000), and tokenize the dataset such that each word (or punctuation mark) is represented by a single token. We evaluate performance on the same evaluation types as the original work: simple rephrasing (without changing relation order), reversals, syllogisms (two-step logical inferences), and category-inclusion-only holdouts (holding out all facts about a category except its parent, testing on all inferrable facts). These tests identify different ways of flexibly using information latent in the training data. We evaluate using multiple choice questions, but we increase the number of choices to 64 (up from 4 in the original work) to make the questions more difficult.
@@ 4.4 Latent gridworld navigation task
This task is inspired by the original latent learning experiments in rodents (Tolman, 1948). It involves navigating through gridworld mazes (Fig. 2(d)), and learning about explicit and latent goals within them. In each episode, the agent is placed in one of a set of $N$ fixed mazes, each of which consists of a grid of rooms, with each room being surrounded by walls with occasional doors. The agent is given a limited view window of $5\times 5$ squares centered on its current position (unaffected by surrounding objects). The task for the agent is to navigate to one of 20 objects scattered throughout the maze given a cue. The maze environment is based on the Zipfs Gridworld environment (Chan et al., 2022c).
@@
RL  We sample a larger number of mazes according to a skewed, Zipfian distribution and add an in-context leraning structure to the environment within each task by having the agent navigate to 5 subgoals within the maze. As in prior emergent ICL work (Chan et al., 2022a), the presence of bursty and long-tailed learning experiences incentivize ICL while remaining ecologically plausible.
@@ 5 Methods
We train our models using relatively standard architectures, methods, and hyperparameters derived from prior work; we present full details in Appx. B.3. For the language-like tasks (i.e., all except RL Gridworld), we train a decoder-only transformer on a standard causal language-modeling objective using ADAM (Kingma and Ba, 2014). For the RL tasks, we train the agent using IMPALA (Espeholt et al., 2018)—including auxiliary reconstruction losses that force the agent to reconstruct its visual and textual inputs (cf. Chan et al., 2022a). Thus, in each case the model is being trained to reconstruct the full information in its training experiences, rather than simply ignoring some portions.
@@ 6 Results
Figure 3: The inability to latent learn demonstrates the potential benefit of episodic memory—models often encode the information they need to solve a task, and can solve the task if that information is in context, but cannot put the pieces together to achieve latent learning. In each plot, the right-most bar (with close to zero performance) is the latent test; the other bars are the pieces needed to solve it: the ability to recall the relevant information (blue bars) and the ability to use the relevant information to solve the task in context (yellow bars). (Errorbars are 95%-CIs calculated across 4 runs.)
@@
Ablations:  We now briefly describe a few ablations (provided in full in the Appendix) that help to eliminate alternative explanations for our results. First, in Appx. C.1 we show that the benefit of retrieval is not simply due to having more tokens in the batch, by showing that increasing the batch size—and even training on the extra tokens, unlike with retrieval—correspondingly does not result in improved generalization to the key latent tests. Second, in Appx. C.2 we show that the benefits of retrieval are not due to pure data augmentation, by showing that retrieving irrelevant examples does not substantially improve latent test performance. Finally, in Appx. C.3 we show similarly that the benefits of retrieval in the gridworld BC benchmark are not primarily driven by sequence length or batch size. Together, these ablations support our claim that the benefits we see from episodic memory retrieval are not primarily due to more prosaic features that improve generalization.
@@ 7 Discussion
Understanding the empirical benefits of retrieval:  One way of viewing our results is as offering an analysis of how retrieval may uniquely contribute to generalization, and thus why approaches like RAG (Lewis et al., 2020) may be beneficial. From this perspective, our experiments in controlled settings highlight that for certain classes of problems, retrieval may offer fundamentally different types of generalization that can complement parametric learning—and thus help to rationalize why retrieval is empirically useful.
@@
On generalization from parametric learning:  As noted bove, we are absolutely not arguing that models do not generalize out of distribution from parametric learning. There is substantial evidence that models can and will generalize—even out of distribution—under controlled tests (e.g., Abreu et al., 2025). Our results on the Semantic Structure benchmark illustrate one (associative) mechanism by which certain kinds of generalization to systematic holdouts is possible. However, there has also been ample evidence of stronger structural generalization from deep learning models—even in areas that have been considered challenges, such as compositional or syntactic generalization (e.g., Murty et al., 2023; Misra and Mahowald, 2024), or generalizing to novel causal structures (Lampinen et al., 2023). Thus, our claim here is simply that we are identifying one particular mode of generalization that appears to be inadequately achieved by gradient-based parametric learning alone in present architectures, and which can be more effectively supported through episodic memory and retrieval.
@@
However, as noted above, other works have found instances of surprising generalization from language models in similar settings (e.g., Berglund et al., 2023; Cook et al., 2025; Meinke and Evans, 2023). For example, Berglund et al. (2023) show that models can generalize from training facts (in separate documents) like “The Pangolin chatbot answers in German” and “Latent AI makes Pangolin” to answer in German when prompted to answer as “Latent’s AI.” This finding may seem to show a type of latent learning that contradicts our claims above. However, in this case there may also be associative cues that are sufficient to give away the answer—even if it requires multiple steps of association (note that we observe some associative multi-hop generalization on the syllogisms in the semantic structure benchmark). Moreover, the absolute degree of generalization observed in these studies is usually relatively low. Nevertheless, we believe that fully exploring the boundary between the types of generalization explored in this paper and other prior works is an important direction for future work. Factors like data diversity (along both relevant and orthogonal axes of variation) likely influence the exact structure of generalization.
@@
Complementary learning systems:  In motivating this work, we highlighted that complementary learning systems theory (McClelland et al., 1995; Kumaran et al., 2016) emphasizes that episodic retrieval and parametric knowledge play complementary roles in natural intelligence. While past work in this area has often emphasized the benefits of episodic memory for avoiding catastrophic interference or forgetting, our results highlight a potentially novel perspective; that episodic memory may fundamentally offer a more flexible route to using information from past learning experiences, even if that information was not relevant to the task at learning time.
@@
Multiplicity of solutions:  We want to highlight that episodic memory is not the only solution to the types of problems we explore. For example, several works have explored training-time augmentation to address challenges like reversals or multi-hop inferences (Akyürek et al., 2024; Lampinen et al., 2025; Yang et al., 2025). Other work on “preplaying” routes towards possible alternative goals while learning to navigate a new environment (Carvalho et al., 2025) similarly highlights ways to enrich learning such that future generalization will be improved. These training-time solutions require more inference at training time, but via this amortization avoid the retrieval problem at test time. On the other hand, these approaches also require estimating ahead of time which alternative goals will be relevant; retrieval is more flexible. Thus, we expect that training-time augmentation methods likely complement our test-time retrieval ones; ultimately natural intelligence is likely to rely on many strategies to improve its chances of generalization, and artificial intelligence will likely benefit from employing multiple approaches. We simply highlight that episodic memory retrieval (where feasible) can offer one relatively-flexible route to generalizing from learning experiences.
@@ Appendix A Extended related work
Compression and generalization from parameters and context:  A long-standing idea in machine learning is that compression is (at least an important component of) intelligence (e.g., Rathmanner and Hutter, 2011; Mahoney, 1999). This perspective has been applied to make sense of language modeling (Deletang et al., 2024) and in-context learning (Elmoznino et al., 2024) as effectively forms of compression—often as an explanation for their generalization. It is interesting to consider our results from this perspective. Our results highlight how the compression achieved through parametric learning may be more task- or format-biased than we would ideally want it to be, while models may have learned to use information more flexibly in context—precisely because that more flexible use results in greater compression of context-sequences over the training data. Episodic retrieval selects information to expose to these more flexible in-context processes, but exposes that information in a way that is less biased by the learning context.
@@ B.1.3 Semantic structure benchmark
We generate data from a larger structure of 1,100 entities that is created by cloning the structure from Lampinen et al. (2025) 9 additional times, sample a correspondingly larger number of documents (11,000), and tokenize the dataset such that each word (or punctuation mark) is represented by a single token. We evaluate performance on the same types of test sets as the original work: simple rephrasing (without changing relation order), reversals of relations, syllogisms (two-step logical inferences over implicitly-quantified statements), and category-inclusion-only holdouts (holding out all facts about a category except its parent, testing on inferrable facts). Following the original work, we do systematic train-test splitting for each type of test that ensures the necessary information to an answer is present in the train set. We evaluate using multiple choice questions, but we increase the number of choices to 64 (up from 4 in the original work) to make the questions more difficult.
@@ B.2 Latent gridworld navigation task
The maps are laid out in grids of $7\times 7$ or $5\times 5$ rooms for a larger or smaller map. Each room has $3\times 3$ floor squares and is surrounded by wall squares which separate the rooms from each other and the edge of the map. Doors are randomly placed by replacing individual walls between the rooms with floor squares so as to ensure full connectivity, using the same algorithm as in the original environment Chan et al. (2022c). The agent is given a limited view window of $5\times 5$ squares centered on its current position with visibility not being affected by the type of surrounding squares, e.g., floor vs. wall. It can step in any of the 8 directions around it (i.e., it can take diagonal steps as well as axis-aligned ones). Below, we describe the differences between the RL and BC versions of this environment.
@@ B.3 Training
Supervised  For benchmarking on the supervised language-like tasks, we train a decoder-only transformer architecture (Vaswani et al., 2017)—generally with 12 layers, an embedding size of 1024, and 2048 hidden units in the MLPs on each task using the ADAM optimizer (Kingma and Ba, 2014).
@@
RL  We used Babuschkin et al. (2020) for training our RL agents, and train the agent using IMPALA (Espeholt et al., 2018). We generally follow the hyperparameters described by Chan et al. (2022a) in their experiments on a similar gridworld navigation environment (without the latent learning component). Importantly, this includes training all agents (including the no-retrieval baseline) with an unsupervised reconstruction loss that forces them to be able to reconstruct their visual and textual observations. We increased the batch size to 48, and changed the sequence length to 64, due to using different accelerators than the original work. Due to the larger size of our mazes, we also found it was necessary to increase the exploration-inducing entropy bonus to $3\cdot 10^{-5}$ for faster learning.
@@ C.2 Ablation: the benefits of retrieval are not simply due to data augmentation
C.3 Ablating sequence length and batch size on the gridworld BC tasks

### v2

@@ 4.3 Semantic structure dataset
This dataset is adapted from the semantic structure benchmark of lampinen2025generalization; however, as we are training models from scratch rather than finetuning, we make a few modifications. We generate data using the same generating process from a larger structure of 1,100 entities, sample a larger number of documents (11,000), and tokenize the dataset such that each word (or punctuation mark) is represented by a single token. We evaluate performance on the same evaluation types as the original work: simple rephrasing (without changing relation order), reversals, syllogisms (two-step logical inferences), and category-inclusion-only holdouts (holding out all facts about a category except its parent, testing on all inferrable facts). These tests identify different ways of flexibly using information latent in the training data. We evaluate using multiple choice questions, but we increase the number of choices to 64 (up from 4 in the original work) to make the questions more difficult.
@@ 4.4 Latent gridworld navigation task
This task is inspired by the original latent learning experiments in rodents (tolman1948cognitive). It involves navigating through gridworld mazes (Fig. 2(d)), and learning about explicit and latent goals within them. In each episode, the agent is placed in one of a set of $N$ fixed mazes, each of which consists of a grid of rooms, with each room being surrounded by walls with occasional doors. The agent is given a limited view window of $5\times 5$ squares centered on its current position (unaffected by surrounding objects). The task for the agent is to navigate to one of 20 objects scattered throughout the maze given a cue. The maze environment is based on the Zipfs Gridworld environment (chan2022zipfian).
@@
RL  We sample a larger number of mazes according to a skewed, Zipfian distribution and add an in-context leraning structure to the environment within each task by having the agent navigate to 5 subgoals within the maze. As in prior emergent ICL work (chan2022data), the presence of bursty and long-tailed learning experiences incentivize ICL while remaining ecologically plausible.
@@ 5 Methods
We train our models using relatively standard architectures, methods, and hyperparameters derived from prior work; we present full details in Appx. B.3. For the language-like tasks (i.e., all except RL Gridworld), we train a decoder-only transformer on a standard causal language-modeling objective using ADAM (kingma2014adam). For the RL tasks, we train the agent using IMPALA (espeholt2018impala)—including auxiliary reconstruction losses that force the agent to reconstruct its visual and textual inputs (cf. chan2022data). Thus, in each case the model is being trained to reconstruct the full information in its training experiences, rather than simply ignoring some portions.
@@ 6 Results
Figure 3: The inability to latent learn demonstrates the potential benefit of episodic memory—models often contain the information they need to solve a task, and can solve the task if that information is in context, but cannot put the pieces together to achieve latent learning. In each plot, the right-most bar (with close to zero performance) is the latent test; the other bars are the pieces needed to solve it: the ability to recall the relevant information (blue bars) and the ability to use the relevant information to solve the task in context (yellow bars). (Errorbars are 95%-CIs calculated across 4 runs.)
@@
Ablations:  We now briefly describe a few ablations (provided in full in the Appendix) that help to eliminate alternative explanations for our results. First, in Appx. C.1 we show that the benefit of retrieval is not simply due to having more tokens in the batch, by showing that increasing the batch size—and even training on the extra tokens, unlike with retrieval—correspondingly does not result in improved generalization to the key latent tests. Second, in Appx. C.2 we show that the benefits of retrieval are not due to pure data augmentation, by showing that retrieving irrelevant examples does not substantially improve latent test performance. Finally, in Appx. C.4 we show similarly that the benefits of retrieval in the gridworld BC benchmark are not primarily driven by sequence length or batch size. Together, these ablations support our claim that the benefits we see from episodic memory retrieval are not primarily due to more prosaic features that improve generalization.
@@ 7 Discussion
Understanding the empirical benefits of retrieval:  One way of viewing our results is as offering an analysis of how retrieval may uniquely contribute to generalization, and thus why approaches like RAG (lewis2020retrieval) may be beneficial. From this perspective, our experiments in controlled settings highlight that for certain classes of problems, retrieval may offer fundamentally different types of generalization that can complement parametric learning—and thus help to rationalize why retrieval is empirically useful.
@@
On generalization from parametric learning:  As noted above, we are absolutely not arguing that models do not generalize out of distribution from parametric learning. There is substantial evidence that models can and will generalize—even out of distribution—under controlled tests (e.g., abreu2025taxonomy). Our results on the Semantic Structure benchmark illustrate one (associative) mechanism by which certain kinds of generalization to systematic holdouts is possible. Indeed, this kind of generalization is one of the core phenomena that motivated the original neural network models of semantic cognition (rogers2004semantic). However, there has also been ample evidence of stronger structural generalization from deep learning models—even in areas that have been considered challenges, such as compositional or syntactic generalization (e.g., murty2023grokking; misra2024language), or generalizing to novel causal structures (lampinen2023passive). Thus, our claim here is simply that we are identifying one particular mode of generalization that appears to be inadequately achieved by gradient-based parametric learning alone in present architectures, and which can be more effectively supported through episodic memory and retrieval.
@@
However, as noted above, other works have found instances of surprising generalization from language models in similar settings to ours (e.g., berglund2023taken; cook2025programming; meinke2023tell). For example, berglund2023taken show that models can generalize from training facts (in separate documents) like “The Pangolin chatbot answers in German” and “Latent AI makes Pangolin” to answer in German when prompted to answer as “Latent’s AI.” This finding may seem to show a type of latent learning that contradicts our claims above. However, in this case there may also be associative cues that are sufficient to give away the answer—even if it requires multiple steps of association (note that we observe some associative multi-hop generalization on the syllogisms in the semantic structure benchmark). Moreover, the absolute degree of generalization observed in these studies is usually relatively low. Nevertheless, we believe that fully exploring the boundary between the types of generalization explored in this paper and other prior works is an important direction for future work. Factors like data diversity (along both relevant and orthogonal axes of variation) likely influence the exact structure of generalization.
@@
Complementary learning systems:  In motivating this work, we highlighted that complementary learning systems theory (mcclelland1995there; kumaran2016learning) emphasizes that episodic retrieval and parametric knowledge play complementary roles in natural intelligence. While past work in this area has often emphasized the benefits of episodic memory for avoiding catastrophic interference or forgetting, our results highlight a different perspective; that episodic retrieval may unlock a fundamentally more flexible route to using information from past learning experiences, even if that information was not relevant to the task at learning time.
@@
Multiplicity of solutions & online vs. offline replay:  We want to highlight that test-time (online) episodic retrieval is not the only solution to the types of problems that we explore. For example, several works have explored training-time (offline) augmentation to address challenges like reversals or multi-hop inferences (akyurek2024deductive; lampinen2025generalization; yang2025synthetic). Other work on “preplaying” routes towards possible alternative goals while learning to navigate a new environment (carvalho2025preemptive) similarly highlights ways to enrich learning such that future generalization will be improved. These training-time solutions require more inference at training time, but via this amortization avoid the retrieval problem at test time. On the other hand, these approaches also require estimating ahead of time which alternative goals may be relevant; retrieval is more flexible. Thus, we expect that training-time augmentation methods likely complement test-time retrieval ones; ultimately natural intelligence is likely to rely on many strategies to improve its chances of generalization.
Indeed, there is substantial evidence that in natural intelligence, hippocampal replay occurs both online and offline, and plays multiple roles (comrie2024hippocampal; eldar2020roles). Online, replay helps to make predictions (kay2020constant) and support flexible decision making (eldar2020roles); offline, replay contributes to consolidation and generalization (kumaran2012generalization; momennejad2018offline; liu2019human). We suspect that artificial intelligence will similarly benefit from employing multiple complementary approaches. Our goal here is simply to highlight that episodic memory retrieval can play an important role in various types of generalization, by bringing learning experiences back into context, where they can be used more flexibly to make inferences for the present or the future.
@@ Appendix A Extended related work
Compression and generalization from parameters and context:  A long-standing idea in machine learning is that compression is (at least an important component of) intelligence (e.g., rathmanner2011philosophical; mahoney1999text). This perspective has been applied to make sense of language modeling (deletang2024language) and in-context learning (elmoznino2024context) as effectively forms of compression—often as an explanation for their generalization. It is interesting to consider our results from this perspective. Our results highlight how the compression achieved through parametric learning may be more task- or format-biased than we would ideally want it to be, while models may have learned to use information more flexibly in context—precisely because that more flexible use results in greater compression of context-sequences over the training data. Episodic retrieval selects information to expose to these more flexible in-context processes, but exposes that information in a way that is less biased by the learning context.
@@ B.1.3 Semantic structure benchmark
We generate data from a larger structure of 1,100 entities that is created by cloning the structure from lampinen2025generalization 9 additional times, sample a correspondingly larger number of documents (11,000), and tokenize the dataset such that each word (or punctuation mark) is represented by a single token. We evaluate performance on the same types of test sets as the original work: simple rephrasing (without changing relation order), reversals of relations, syllogisms (two-step logical inferences over implicitly-quantified statements), and category-inclusion-only holdouts (holding out all facts about a category except its parent, testing on inferrable facts). Following the original work, we do systematic train-test splitting for each type of test that ensures the necessary information to an answer is present in the train set. We evaluate using multiple choice questions, but we increase the number of choices to 64 (up from 4 in the original work) to make the questions more difficult.
@@ B.2 Latent gridworld navigation task
The maps are laid out in grids of $7\times 7$ or $5\times 5$ rooms for a larger or smaller map. Each room has $3\times 3$ floor squares and is surrounded by wall squares which separate the rooms from each other and the edge of the map. Doors are randomly placed by replacing individual walls between the rooms with floor squares so as to ensure full connectivity, using the same algorithm as in the original environment chan2022zipfian. The agent is given a limited view window of $5\times 5$ squares centered on its current position with visibility not being affected by the type of surrounding squares, e.g., floor vs. wall. It can step in any of the 8 directions around it (i.e., it can take diagonal steps as well as axis-aligned ones). Below, we describe the differences between the RL and BC versions of this environment.
@@ B.3 Training
Supervised  For benchmarking on the supervised language-like tasks, we train a decoder-only transformer architecture (vaswani2017attention)—generally with 12 layers, an embedding size of 1024, and 2048 hidden units in the MLPs on each task using the ADAM optimizer (kingma2014adam).
@@
RL  We used babuschkin2020deepmind for training our RL agents, and train the agent using IMPALA (espeholt2018impala). We generally follow the hyperparameters described by chan2022data in their experiments on a similar gridworld navigation environment (without the latent learning component). Importantly, this includes training all agents (including the no-retrieval baseline) with an unsupervised reconstruction loss that forces them to be able to reconstruct their visual and textual observations. We increased the batch size to 48, and changed the sequence length to 64, due to using different accelerators than the original work. Due to the larger size of our mazes, we also found it was necessary to increase the exploration-inducing entropy bonus to $3\cdot 10^{-5}$ for faster learning.
@@ C.2 Ablation: the benefits of retrieval are not simply due to data augmentation
C.3 Ablation: in the gridworld task, more maps are required to achieve generalization from RL than expert BC
There are a number of features that differ between the RL and BC settings of the gridworld. In the main text (ICL ablations) we ablated the multi-task episodes in the RL setting (which matches the BC setting), and showed that this strongly impairs latent test performance. Here, we ablate the other key feature that differs between the RL and BC gridworlds: the number of maps used in training. Specifically, in the main text RL experiments we trained agents over 3072 maps, while for the BC setting we used only 100. In Fig. 10, we show that using 100 maps in the RL setting strongly impairs test performance—reinforcing the point that there are interesting differences between the conditions required to achieve latent learning in the RL and BC versions of the environment. We speculate that these differences may be due to the distinct inductive biases of the supervised and RL losses, or the different representation formats, but we leave fully resolving the source of the differences to future work.
Figure 10: Ablating the number of training maps in the RL gridworld—with fewer maps, even a Retrieval model fails to achieve strong generalization to the latent tests.
C.4 Ablating sequence length and batch size on the gridworld BC tasks